Find two consecutive integers such that 4 times the larger is 28 more than 3 times the smaller.
Answer by ikleyn(53213)

.

The basic equation is


    4*(n+1) = 3n + 28.


    4n + 4 = 3n + 28

    4n - 3n = 28 - 4

     n      = 24.


ANSWER.  The numbers are  24 and 25.

Solved.